OpenStudy (anonymous):

If f(x) = x^2 + 1 and g(x) = 3x + 1, find 2f(1) + 3g(4)

OpenStudy (anonymous):

well,first find f(1)...you must only substitute 1 instead of x here:x^2+1 and for g(4) u must substitute 4 instead of x : 3x+ 1 so, f(1) = 1+1=2 g(4)=12+1=13 now can u find 2f(1)+3g(4) ?

OpenStudy (anonymous):

i'm lost.. I'm not good at math at all...